OpenStudy (anonymous):

Introduction to famous curves :- Rose curve I thought it a good idea if i posted this topic each week maybe , I'll choose a famous cool curve each time to talk about, ready to answer any question about it, just feel free to ask and maybe share cool ideas :) i hope you to have fun with it !

OpenStudy (anonymous):

A curve which has the shape of a petalled flower. This curve was named rhodonea by the Italian mathematician Guido Grandi between 1723 and 1728 because it resembles a rose (MacTutor Archive). The polar equation of the rose is \(r=a\sin(n\theta)\) or \(r=acos(n\theta)\).

OpenStudy (anonymous):

I'll start with \(r=a\sin(n\theta)\) \(a\) affects on expanding and sketching i'll show you n effects only when \(n\) is \(0\) the graph is a point \( \sin 0=0\) . case 1:- the standard graph when \(n=1\) http://prntscr.com/5vq2z2 case 2:- when \(|n|<1\) note when its negative the its reflective around \(x\) axis so when \(n=1/m\) and \(m\) is even no change happens . http://prntscr.com/5vq3xu when \(n=1/m\) and \(m\) is odd http://prntscr.com/5vq88v case 3:- when \(|n|>1\) the graph would be just like petal and u'll note \(n\) is the petals numbers , when \(n\) is negative odd is a reflexive of the graph when \(n\) is positive around \(x\) axis http://prntscr.com/5vqfi7 when \(n\) is even then the number of petals is \(2n\) http://prntscr.com/5vqke2 FOR MORE FUN https://www.desmos.com/calculator/tyuk4he0e3

OpenStudy (anonymous):

i'll continue on this later since i dint add the spiral case when n is too small

ganeshie8 (ganeshie8):

that desmos link is amazing xD

OpenStudy (anonymous):

haha yeah

OpenStudy (kainui):

https://www.desmos.com/calculator/n5ussfeanz

Nnesha (nnesha):

hmm waoo thatz cool ^^^^^^looks amazing

OpenStudy (anonymous):

aww that so nice :O